Please help me solve this inequality:
2x^2 + x >= 21

Answer by tutorcecilia(2144)

2x^2 + x >= 21 [set the inequality equal to zero]
2x^2 + x -21>= 21-21
2x^2 + x -21>= 0 [Factor]
(2x+7)(x-3)>= 0
(2x+7)(x-3)>= 0 [Set each factor >=0]
2x+7>=0 [Solve for x]
2x>=-7/2
.
x-3>=0
x>=3
.
So, x>=-7/2 or x>=3
.
Check each answer by plugging each value of x back into the original inequality to see if the statement is true.